### v1

@@ Abstract
Orthogonal nonnegative matrix factorization (ONMF) has become a standard approach for clustering. As far as we know, most works on ONMF rely on the Frobenius norm to assess the quality of the approximation. This paper presents a new model and algorithm for ONMF that minimizes the Kullback-Leibler (KL) divergence. As opposed to the Frobenius norm which assumes Gaussian noise, the KL divergence is the maximum likelihood estimator for Poisson-distributed data, which can model better vectors of word counts in document data sets and photo counting processes in imaging. We have developed an algorithm based on alternating optimization, KL-ONMF, and show that it performs favorably with the Frobenius-norm based ONMF for document classification and hyperspectral image unmixing.
@@ 1 Introduction
where $\|X\|_{F}^{2}=\sum_{i,j}X_{i,j}^{2}$ is the squared Frobenius norm. The only exceptions we were able to find are the papers [11, 8] in which the authors consider the KL and Bregman divergences nut they used a regularization term to promote orthogonality, leading to soft clusterings; in this paper we impose orthogonality as a hard constraint. The underlying assumption when using the Frobenius norm is that the noise follows a Gaussian distribution. However, in situations when Gaussian noise is not meaningful, other objective functions should be used. In particular, and this will be the focus of this paper, if the entries of the data matrix follow a Poisson distribution of parameter given by $WH$, then one should minimize the Kullback-Leibler (KL) divergence between $X$ and $WH$, which is defined as
@@
The matrix $X\in\mathbb{R}^{m\times n}$ is a real $m$-by-$n$ matrix, its $j$th column is denoted $X(:,j)$, its $i$th row by $X(i,:)$, its entry at position $(i,j)$ by $X_{i,j}$, its transpose by $X^{\top}$, and its number of non-zero entries by $\operatorname{nnz}(X)$. Given sets of indices $\mathcal{I}$ and $\mathcal{K}$, $X(\mathcal{I},\mathcal{J})$ denotes the submatrix of $X$ with row (resp. column) indices in $\mathcal{I}$ (resp. $\mathcal{J}$). The vectors $e$ is the vector of all ones of appropriate dimension, and $I_{r}$ is the identity matrix of dimension $r$.
@@ 3.2 Update $H$ when $W$ is fixed
$f(\alpha)=D_{\text{KL}}(X(:,j),\alpha W(:,k))$ $=\alpha e^{\top}W(:,k)-\sum_{i=1}^{m}X(i,j)\log(\alpha W(i,k))$
$=\alpha e^{\top}W(:,k)-e^{\top}X(:,j)\log\alpha-\sum_{i=1}^{m}X(i,j)\log W(i,k).$
@@
The entry of the column of $H(:,j)$ that should be set to a positive value is the one that minimizes the quantity $KL(X(:,j),H_{k,j}^{*}W(:,k))$, hence it is the entry corresponding to the index $k$ such that $\sum\limits_{i=1}^{m}X(i,j)\log\left(\frac{W(i,k)}{e^{\top}W(:,k)}\right)$ is maximized. That is, to select the column of $W$ that best approximates the column $X(:,j)$ in the KL divergence, we pick the column $W(:,k)$ maximizing [^2] Recall that, by convention, $00\log 0=$ 0, as KL$(0,y)=y$. $\sum\limits_{i=1}^{m}X(i,j)\log\left(\frac{W(i,k)}{e^{\top}W(:,k)}\right)$. Numerically, we can construct $W_{n}$ by normalizing the columns of $W$ to have unit $\ell_{1}$, and then compute $A:=\log(W_{n}+\epsilon)^{\top}X$, where the $\log$ is taken component-wise: the locations of the non-zero entries of the optimal $H$ are given by the maximum of each column of $A$.
@@ 4 Numerical Experiments
In this section, we compare the performance of Fro-ONMF (Algorithm 1) and KL-ONMF (Algorithm 2) for clustering documents (Section 4.1) and pixels in hyperspectral images (Section 4.2). The code is available from https://gitlab.com/ngillis/kl-onmf, and can be used to rerun all experiments presented below. All experiments were run on a LAPTOP Intel(R) Core(TM) i7-8850H CPU @ 2.60GHz 16,0Go RAM.
@@
There are many ways to initialize ONMF algorithms, as they are for $k$-means. To simplify the presentation, we use in this paper the approach proposed in [6]. It initializes $W$ with the successive nonnegative projection algorithm (SNPA) [5] that identifies a subset of $r$ columns of $X$ that represent well-spread data points in the data set.

### v2

@@ Abstract
Orthogonal nonnegative matrix factorization (ONMF) has become a standard approach for clustering. As far as we know, most works on ONMF rely on the Frobenius norm to assess the quality of the approximation. This paper presents a new model and algorithm for ONMF that minimizes the Kullback-Leibler (KL) divergence. As opposed to the Frobenius norm which assumes Gaussian noise, the KL divergence is the maximum likelihood estimator for Poisson-distributed data, which can model better sparse vectors of word counts in document data sets and photo counting processes in imaging. We develop an algorithm based on alternating optimization, KL-ONMF, and show that it performs favorably with the Frobenius-norm based ONMF for document classification and hyperspectral image unmixing.
@@ 1 Introduction
where $\|X\|_{F}^{2}=\sum_{i,j}X_{i,j}^{2}$ is the squared Frobenius norm. The only exceptions we were able to find are the papers [11, 8] in which the authors consider the KL and Bregman divergences but they used a regularization term to promote orthogonality, leading to soft clusterings; in this paper we impose orthogonality as a hard constraint. The underlying assumption when using the Frobenius norm is that the noise follows a Gaussian distribution. However, in situations when Gaussian noise is not meaningful, other objective functions should be used. In particular, and this will be the focus of this paper, if the entries of the data matrix follow a Poisson distribution of parameter given by $WH$, then one should minimize the Kullback-Leibler (KL) divergence between $X$ and $WH$, which is defined as
@@
The matrix $X\in\mathbb{R}^{m\times n}$ is a real $m$-by-$n$ matrix, its $j$th column is denoted $X(:,j)$, its $i$th row by $X(i,:)$, its entry at position $(i,j)$ by $X_{i,j}$, its transpose by $X^{\top}$, and its number of non-zero entries by $\operatorname{nnz}(X)$. Given sets of indices $\mathcal{I}$ and $\mathcal{J}$, $X(\mathcal{I},\mathcal{J})$ denotes the submatrix of $X$ with row (resp. column) indices in $\mathcal{I}$ (resp. $\mathcal{J}$). The vectors $e$ is the vector of all ones of appropriate dimension, and $I_{r}$ is the identity matrix of dimension $r$.
@@ 3.2 Update $H$ when $W$ is fixed
$f(\alpha)=D_{\text{KL}}(X(:,j),\alpha W(:,k))$ $=\alpha e^{\top}W(:,k)-\sum_{i=1}^{m}X(i,j)\log(\alpha W(i,k))\vspace{-0.2cm}$
$=\alpha e^{\top}W(:,k)-e^{\top}X(:,j)\log\alpha-\sum_{i=1}^{m}X(i,j)\log W(i,k).\vspace{-0.2cm}$
@@
The entry of the column of $H(:,j)$ that should be set to a positive value is the one that minimizes the quantity $KL(X(:,j),H_{k,j}^{*}W(:,k))$, hence it is the entry corresponding to the index $k$ such that $\sum\limits_{i=1}^{m}X(i,j)\log\left(\frac{W(i,k)}{e^{\top}W(:,k)}\right)$ is maximized. That is, to select the column of $W$ that best approximates the column $X(:,j)$ in the KL divergence, we pick the column $W(:,k)$ maximizing [^2] Recall that, by convention, $00\log 0=$ 0, as KL$(0,y)=y$. $\sum\limits_{i=1}^{m}X(i,j)\log\left(\frac{W(i,k)}{e^{\top}W(:,k)}\right)$. Numerically, we can construct $W_{n}$ by normalizing the columns of $W$ to have unit $\ell_{1}$, and then compute $A:=\log(W_{n}+\epsilon)^{\top}X$, where $\epsilon$ is a small constant to avoid numerical errors, and the $\log$ is taken component-wise: the locations of the non-zero entries of the optimal $H$ are given by the maximum of each column of $A$.
@@ 4 Numerical Experiments
In this section, we compare the performance of Fro-ONMF (Algorithm 1) and KL-ONMF (Algorithm 2) for clustering documents (Section 4.1) and pixels in hyperspectral images (Section 4.2). The code in MATLAB is available from https://gitlab.com/ngillis/kl-onmf, and can be used to rerun all experiments presented below. All experiments were run on a LAPTOP Intel(R) Core(TM) i7-8850H CPU @ 2.60GHz 16,0Go RAM.
@@
There are many ways to initialize ONMF algorithms, as they are for $k$-means. To simplify the presentation, we use the approach proposed in [6]. It initializes $W$ with the successive nonnegative projection algorithm (SNPA) [5] that identifies a subset of $r$ columns of $X$ that represent well-spread data points in the data set.